# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v3

@@ 5.3 Results of LLMs with Formula Retriever
Model zero-shot few-shot GLM4 65.32 62.47    + formula retriever 70.31 65.80 GPT-4-turbo 70.07 71.50    + formula retriever 68.17 67.00
@@
The results of LLMs utilizing the formula retriever are shown in Table 5. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on GLM4, with score increments of 4.99 and 3.33 with zero-shot and few-shot, respectively, on the id test. However, we observed a performance decline with GPT-4-turbo. Specifically, we found that the top 5 retrieved formulas often included irrelevant ones, as the number of formulas required varies for each problem. The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in utilizing a formula database.
@@ 5.4 Results of Supervised Fine-tuned Models
Model Size id test ood test Avg. Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
@@
Table 6 shows the results for the supervised fine-tuned models, with and without CoT-SFT and DA, which were detailed in Section 4.2.4. In most settings, both models achieved higher scores on the id test than the ood test, yet they still exhibited considerable performance on the ood test. This indicates that 1) the ood formulas indeed impacted model performance and 2) the models still demonstrate generalizability. We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future works.
It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4-turbo (though such a comparison may not be entirely fair). This indicated that the incorporation of CoT-SFT and the use of calculators could significantly enhance the reasoning capabilities of small models. Our findings revealed that focusing on reasoning with CoT while delegating numerical calculation to a calculator could enhance the performance of small models, given their limited calculating capability. The assistance of LLMs for data augmentation could also enhance smaller models’ reasoning capability. This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
@@ 6 Conclusion and Limitations
We have also translated the dataset into English unitizing LLMs. However, we have not yet accurately assessed the quality of the translated dataset. At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset. Additionally, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
Acknowledgments and Disclosure of Funding
This work was supported by the CIPSC-SMP-Zhipu.AI Large Model Cross-Disciplinary Fund. We are grateful to Chao Li for his suggestions and efforts in the annotation.
@@ A.1.3 Semantic-based Merging for Formula Database Construction
A.2 Case Study and Error Analysis
(a) An error case caused by wrong formulas.
(b) An error case caused by wrong calculation.
Figure 2: Error cases.
A.3 Experiments
A.3.1 Data Augmentation (DA) for FormulaReasoning
@@
We found that mixing the newly generated data into the original training set did not always bring positive improvement, perhaps because the newly generated data has not undergone manual review. We found that randomly selecting a small portion of the newly generated data can enable the model to have performance improvement. We set several different mixing ratios selected from $\{5\%,10\%,15\%,20\%,2\%,30\%,35\%,40\%\}$. We fine-tuned the ChatGLM-6B-base using the augmented data set. After training for a fixed number of steps (150k and 200k), we selected the checkpoints with the smallest loss among models of different mixing ratios.
A.4 Implementation Details
We accessed to GPT-4-turbo, GPT-3.5-turbo [^5] https://platform.openai.com/docs, GLM4666https://open.bigmodel.cn/, and Qwen-max [^7] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For ChatGLM3-6B, we fine-tuned with LoRA Hu et al. (2021) with r=8, alpha=32 and learning rate of 5e-5, batch size of 1. The max input length and output length are both set to 512. We utilized nucleus sampling with top_p=0.8 for generation. In the case of CoT-SFT, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training mT5-base, mT5-large, Qwen-1.8B, ChatGLM-6B models requires 6, 12, 12 and 24 hours respectively.
A.4.1 Zero-shot and Few-shot Prompts
@@ A.4.2 Formula Retriever
A.4.3 Few-shot Experiments on the LLMs of Chat Versions
Model Size id test ood test Avg. zero-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Llama3-instruct 8B 22.66 17.98 20.41 Llama2-chat 7B 0.00 0.26 0.13 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Llama3-base 8B 12.81 8.87 10.91 Llama2-base 7B 1.43 0.26 0.87 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 few-shot CoTs with LLMs of chat/instruct versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Llama3-instruct 8B 6.90 6.16 6.54 Llama2-chat 7B 1.97 1.00 1.50 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03
Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaReasoning.
In this experiment, we compared the performance of the same version of the model under zero-shot and few-shot settings. Results are shown in Table 7. For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test. Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease from base to chat versions.
A.4.4 Prompts for LLMs with Formula Retriever

### v4

@@ 5.3 Results of LLMs with Formula Retriever
Model zero-shot few-shot GLM-4-flash 71.50 61.76    + formula retriever 70.55 62.95 Qwen2.5-7B 42.04 59.62    + formula retriever 52.96 63.66
The results of LLMs utilizing the formula retriever are shown in Table 5. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on Qwen2.5-7B, with score increments of 10.92 and 4.04 with zero-shot and few-shot, respectively, on the id test. However, we found that the performance was essentially on par on the GLM-4-flash. Specifically, we found that the top 5 retrieved formulas often included irrelevant ones, as the number of formulas required varies for each problem. The presence of these extraneous formulas affected the model’s performance, indicating that there is considerable room for further research in retrieving from a formula database.
@@ 5.4 Results of Supervised Fine-tuned Models
Model Size id test ood test Avg. Qwen-1.8B 1.8B 55.91 44.58 50.25    + DA 56.16 45.32 50.74    + CoT-SFT 73.65 74.38 74.00 ChatGLM3-6B 6B 52.95 40.64 47.02    + DA 53.44 45.32 49.53    + CoT-SFT 74.63 73.89 74.23
Table 6 shows the results for the supervised fine-tuned models, with and without CoT-SFT and DA, which were detailed in Section 4.2.4. In most settings, both models achieved higher scores on the id test than the ood test, yet they still exhibited considerable performance on the ood test. This indicates that 1) the ood formulas indeed challenged model performance and 2) the models still demonstrated a certain level of generalizability. We hope that the division of id test and ood test will be helpful for assessing the generalization ability of fine-tuned models in future work.
It was noteworthy that with CoT-SFT, Qwen-1.8B and ChatGLM3-6B, with a mere parameter count of 1.8B and 6B, respectively, achieved performance comparable to GPT-4o (though such a comparison may not be entirely fair). This indicated that the incorporation of CoT-SFT and the use of calculators could significantly enhance the reasoning capabilities of small models. Our findings revealed that focusing on reasoning with CoT while delegating numerical calculation to a calculator could enhance the performance of small models, given their limited calculating capability. The assistance of LLMs for data augmentation could also enhance smaller models’ reasoning capability. This finding provides valuable insights for future deployment of numerical reasoning systems with small models.
5.5 Case Study and Error Analysis
(a) An error case caused by wrong formulas.
(b) An error case caused by wrong calculation.
Figure 2: Error cases.
@@ 6 Conclusion and Limitations
In the future work, we plan to utilize the formula knowledge from FormulaReasoning to improve the numerical reasoning capabilities of LLMs. Possible approaches include enhancing reasoning abilities through knowledge-driven methods, preference learning methods based on formula feedback.
One limitation of this work is that our evaluation results reported in the paper were obtained from the original Chinese version of FormulaReasoning. We have employed a combination of LLM-based translation and manual review to release an English version of FormulaReasoning. Currently, we provide a preview English version in our GitHub repository, and we will release the official English version of FormulaReasoning after completing the manual review process. Another limitation is that, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people which would benefit evaluation efforts, it is still possible to explore formula-based question answering data in other domains.
@@ A.1.3 Semantic-based Merging for Formula Database Construction
A.2 Experiments
A.2.1 Data Augmentation (DA) for FormulaReasoning
@@
We found that mixing the newly generated data into the original training set did not always bring positive improvement, perhaps because the newly generated data has not undergone manual review. We found that randomly selecting a small portion of the newly generated data can enable the model to have performance improvement. We set several different mixing ratios selected from $\{5\%,10\%,15\%,20\%,2\%,30\%,35\%,40\%\}$. We fine-tuned each model using the augmented data set. After training for a fixed number of steps (150k and 200k), we selected the checkpoints with the smallest loss among models of different mixing ratios.
A.2.2 Zero-shot and Few-shot Prompts
@@ A.2.3 Formula Retriever
A.2.4 Comparison of CoT and PoT Prompts
Table 7: Results of LLMs with zero-shot and few-shot chain of thought (CoT) and program of thought (PoT).
Model zero-shot few-shot id test ood test Avg. id test ood test Avg. GPT-4o (CoT) 77.20 72.38 74.88 76.01 73.66 74.88 GPT-4o (PoT) 80.76 73.91 77.46 81.47 82.61 82.02 GLM-4-plus (CoT) 84.32 81.07 82.76 82.90 85.68 84.24 GLM-4-plus (PoT) 84.08 78.51 81.40 86.70 84.91 85.84 Human 93.49 90.47 92.03 93.49 90.47 92.03
Results are shown in Table 7. In the PoT approach, we utilized a Python interpreter to execute the code and obtain the final results. We found that the performance comparison between CoT and PoT varies across models. GPT-4o consistently demonstrated superior performance with PoT across all settings, achieving improvements of 2.58 points on average in the zero-shot setting and 7.14 points on average in the few-shot setting. In contrast, GLM-4-plus showed an average decline of 1.36 points in the zero-shot setting but showed an average improvement of 1.60 points in the few-shot setting. The finding might be related to the code capabilities of the models.
A.2.5 Prompts for LLMs with Formula Retriever